OpenStudy (anonymous):

Jose has 9 more comic books than Robin. Robin has 5 more comic books than Lee. a. If Jose has 23 comic books, how many comic books does Robin have? b. How many comic books does Lee have?

OpenStudy (anonymous):

\(j = 9 + r\) \(r = 5 + \ell\) \(j = 23\)

OpenStudy (anonymous):

\(r = j - 9 = 23 - 9 = 14\) (robin)

OpenStudy (anonymous):

\(\ell = r - 5 = 15 - 5 = 9\) (lee)